Two trains leave from the same station on the same tracks. The train that is traveling east is going 8mph faster than the train that is going west. After 2 hours and 30 minutes they are 194 miles apart.how fast were the two trains traveling?

Answer by rfer(16322)

194/2.5=77.6 mi in 1 hr
x+x+8=77.6
2x=69.6
x=34.8 mph
x+8=42.8 mph